OpenStudy (anonymous):

The ratio of girls to boys who participated in the quiz bowl was 7:5. There were 42 girls in the competition. How many boys participated? A) 28 B) 30 C) 35 D) 49

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

You can set up a proportion: \(\dfrac{7}{5} = \dfrac{42}{x}\) Cross multiply: \(7 \times x = 7x\) \(5 \times 42 = 210\) \(7x = 210\) Now solve for x. Divide 7 to both sideds, what's 210 / 7?

OpenStudy (anonymous):

30 :)

OpenStudy (anonymous):

Yep, so that's your answer.

OpenStudy (anonymous):

Check: 7 / 5 = 1.4 42 / 30 = 1.4 So 30 is your answer.

OpenStudy (anonymous):

Thank you!!!:)

OpenStudy (anonymous):

No problem :D